OpenStudy (anonymous):

Write an equation in point-slope form of the line through point J(4,1) with slope -4.

OpenStudy (anonymous):

y-y1 = m(x-x1) where m = slope = -4 and (x1,y1) = 4,1)

OpenStudy (kesumonu):

pardon me it should have been 4x+y-17=0

OpenStudy (anonymous):

yes..Good!!

OpenStudy (anonymous):

Thanks to both of you!!!!!! :D